A block of ice with mass 1.00kg is initially at rest on a frictionless, horizontal surface. A worker then applies a horizontal force F to it. A a result, the block moves along the x-axis such that its position as a function of time is given by x(t) = 0.100(m/s2) t2 + 0.0100(m/s3) t3. Calculate the work done by the force during the first 1.00s of the motion. 



1
Expert's answer
2021-06-03T18:26:41-0400

m= 1 kg

"x(t) = 0.100 t^2 + 0.0100t^3"

The acceleration of block can be calculated by differentiating position function twice,

"x’(t) = 0.200t + 0.0300t^2 \\

x’’(t) = a = 0.200 + 0.0600t"

Acceleration at t=1 s is:

"a = 0.200 + 0.0600 \times 1 = 0.2600 \;m/s^2"

Force on block can be calculated as:

"F=ma = 1 \;kg \times 0.26 \;m/s^2 = 0.26 \;kg \; m/s^2"

The position of block is at t=1 s.

"x(1) = 0.100 \times 1^2 + 0.0100 \times 1^3 = 0.11 \;m"

Work done by block is:

"W=F \times x \\

= 0.26\times 0.11 = 0.0286 \;J"

Thus, work done is 0.0286 J.